# Efficient Monte-Carlo sampling of metastable systems using non-local collective variable updates

### v2

@@ II.2.4 RATTLE dynamics and work
where $({\mathcal{G}}^{k})$ and $({\mathcal{G}}^{k+1/2})$ are independent sequences of i.i.d. centered Gaussian random vectors with covariance matrix $\mathrm{Id}_{d}$. The Lagrange multiplier $\lambda^{k+1/2}$ (resp. $\lambda^{k+3/4}$) is chosen so that the constraint $(C_{q})$ (resp. $(C_{p})$) is satisfied. For $\lambda^{k+3/4}$, this leads to a linear equation with an analytical solution, whereas the solution for $\lambda^{k+1/2}$ will in general be obtained numerically using Newton’s methods. For further precisions on this point see Remark˜B.3 in the appendix. Note that there is an equivalent formulation of (II.17) and (II.19) that is more suitable for numerical implementation, which we provide in Appendix˜A for completeness. The matrices $\gamma_{P}(q^{k})=P_{M}(q)\gamma P_{M}(q)^{\top}$ and $\sigma_{P}(q^{k})=P_{M}(q)\sigma$ satisfy the fluctuation-dissipation identity
@@
Assuming that the position and velocity schedules are compatible in the continuous-time limit $\Delta t\to 0$, i.e. $v_{z}(t)=\dot{z}(t)$, Eqs.˜II.17, II.18 and II.19 are discretizations of Langevin dynamics with constraints $(\xi(q^{k}),v_{\xi}(q^{k},p^{k}))=(z(t_{k}),v_{z}(t_{k}))$ see Ref. lelievre_langevin_2012.
@@
for $k=0\ldots K_{T}-1$, where $\widetilde{H}$ denotes the Hamiltonian that is modified by the Fixman term (II.16), namely $\widetilde{H}(q,p)=\tilde{V}(q)+\frac{1}{2}p^{\top}M^{-1}p$}.
For $\gamma=0$, the RATTLE dynamics (II.17)-(II.19) become deterministic (driven Hamiltonian dynamics) and the only remaining source of randomness in Step 2 is the initialization of the momentum $p^{0}$. Furthermore, since in this case $p^{k+1/4}=p^{k}$ and $p^{k+3/4}=p^{k+1}$, the work simplifies to a telescopic sum and just depends on the endpoints of the path:
@@
Remark II.3.
When the Gram tensor $G_{M}(q)$ does not depend on $q$ (an example is given in Section˜III.3), the Fixman term just adds an overall shift to the energy $V$ and need not be considered. In this case, the requirement on the initial and final velocity can also be relaxed to $v_{z}(t_{0})=v_{z}(t_{K_{T}})$ and can be non-zero, assuming that the initial momentum $p^{0}$ is still sampled from the appropriate subspace (see Remarks˜B.2 and B.10). This is true in particular for the case of a linear CV, which, after a change of basis, can always be seen as representing a subset of $\ell$ degrees of freedom. This is the standard setting treated in previous worksathenesComputationChemicalPotential2002; nilmeier2011nonequilibrium; chen_enhanced_2015; chen2015generalized; neal2005taking, where some degrees of freedom are driven through a non-equilibrium protocol.
The problem of constructing a steered path can be understood as transporting samples from the constrained target measure at the initial CV value $Z_{n}$ to the final one $\widetilde{Z}_{n+1}$. Our method is based on a stochastic differential equation (with the auxiliary variable of the momentum), but other dynamics are possible, notably deterministic transformationschen_enhanced_2015 or a combination of stochastic and deterministic steps, where the underlying SDE is modified by a deterministic ‘escorting’ termvaikuntanathan_escorted_2011. The latter can also be understood as an instance of a stochastic normalizing flowcaselle_stochastic_2022. These approaches are not easy to adapt to our context of a nonlinear CV, however, since they require to transport samples supported on the submanifolds $\Sigma_{z}$.
@@ II.3 Reversibility of the algorithm
We prove this statement in Appendix˜B. The basic ingredient of the proof is to introduce a backward version of the dynamics (associated to the forward dynamics) and use a version of the Jarzynski–Crooks equalityCrooks99.
As a consequence of Theorem˜II.4, the MCMC algorithm admits $\nu$ as an invariant measure and, up to checking irreducibility which has to be done on a case-by-case basis, empirical averages over trajectories $(Q_{n})_{n\geq 0}$ converge to averages with respect to $\nu$.
Remark II.5.
@@ II.4 A normalized version of the algorithm
The overall algorithm depends on many parameters: $\Delta t$, $\gamma$, $\sigma$, $M$, and the choice of the schedule $z(t)$ and $v_{z}(t)$ and transition time $T$. In order to reduce the number of parameters, we will use the following normalized version in the numerical examples: Assuming that the mass matrix $M=M\,\mathrm{Id}$ is isotropic, we show in Appendix˜C that the overall dynamics only depends on the two parameters
@@
For the parameter choice $\alpha_{1}=1$, the steered dynamics (II.17)-(II.19) reduce to overdamped Langevin dynamics (see Appendix˜D). If $\xi$ is linear in addition, so that after a suitable change of basis the CV can be identified with the first $\ell$ degrees of freedom, the overall algorithm is equivalent to the ‘asymmetric’ algorithm presented in our earlier work SCHONLE2025113806 (up to a slight change of the definition of the steered path: here we first move the CV and then the remaining degrees of freedom). Therein, we also provided the continuous-time limit for this version of the algorithm together with the corresponding stochastic differential equation (SDE). For the choice of $\alpha_{1}=0$, the steered dynamics correspond to the Verlet scheme (Leapfrog dynamics) in the orthogonal degrees of freedom and are therefore a discretization of Hamiltonian dynamics with a time-dependent Hamiltonian. Note that in this case, the continuous-time limit of the dynamics $\Delta t\to 0$ simply corresponds to taking $\alpha_{2}\to 0$ and the acceptance of the move in Step 3 is governed only by the (normalized) physical transition time $T/\sqrt{\beta M}=K_{T}\Delta t/\sqrt{\beta M}=\mathrm{ceil}\left(|\widetilde{Z}_{n+1}-Z_{n}|/\widetilde{v}\right)\sqrt{\alpha_{2}}$ (see Fig.˜14 in the appendix for a numerical illustration).
@@ III Numerical examples
In this section, we apply the proposed CV-guided sampler to examples of increasing complexity, demonstrating in particular the advantage of deterministic underdamped steering dynamics compared to the previously considered overdamped version. We consider four model systems of increasing complexity: a theoretical model-system dubbed ‘Gaussian tunnel’ (Section˜III.1), the one-dimensional $\phi^{4}$ model from statistical physics (Section˜III.2), a dimer in a solvent (Section˜III.3), and finally a 9-bead polymer in a solvent (Section˜III.4).
@@ III.1 Gaussian tunnel
We first consider a simple model with a linear CV as already introduced in previous workSCHONLE2025113806. It is a toy model in $d=20$ dimensions with a designated one-dimensional variable $z$ serving as the CV and d1d-1 additional coordinates $(x^{\perp}_{i})_{1\leq i\leq d-1}$. The distribution of $z$ is given by a Gaussian mixture and the conditional distribution of the remaining coordinates is a Gaussian with a mean depending on $z$. More precisely,
@@
Running a rough optimization through a grid search reveals that the optimum performance is obtained for $\alpha_{1}=0$, which corresponds to a deterministic steered dynamics (see Fig.˜13 of Section˜H.1). We compare the deterministic ($\alpha_{1}=0$) and the overdamped ($\alpha_{1}=1$) steered dynamics in Fig.˜1 using the two performance measures, which show consistent results. The deterministic dynamics offer a large performance increase over the overdamped dynamics with a difference of around two orders of magnitude. In both cases, we observed that for too large values of $\alpha_{2}$ (a too coarse time discretization), the dynamics become unstable and the trajectories diverge, leading to zero acceptance of the overall move. The optimal value for $\alpha_{2}$ is then just below this threshold. Typical driven transition paths are shown for the optimum of the overdamped and the deterministic case in Fig.˜2. To verify the unbiasedness of the algorithm at the optimum, we show the marginals of $z$ and $x_{0}^{\perp}$ in Fig.˜3. They indeed agree with the target measure and show that the algorithm corrects for the biased proposal distribution in CV space.
@@ III.2 $\phi^{4}$ Model
As a second example we consider a field model from statistical physics, the one-dimensional $\phi^{4}$ model (also considered previouslySCHONLE2025113806). It consists of $N$ degrees of freedom $\{\phi_{i}\}_{i=1}^{N}$ on a one-dimensional lattice and can be seen as a continuous version of the Ising model for ferromagnetism in solids. The target distribution is the Boltzmann–Gibbs probability distribution $\nu({\rm d}\phi)$ with the energy (using the simplified version of the model considered in Ref. gabrieAdaptiveMonteCarlo2022)
@@
We use the magnetization $\overline{\phi}$ as the collective variable, which is a natural order parameter of the system. We run the algorithm on the system in a wavelet basis (see Appendix˜E for details) since this order parameter then naturally emerges as the first component of the system. To fix the proposal sampler in CV space, we first ran a short simulation using the Metropolis Adjusted Langevin Algorithm (MALA) RDF78; roberts1996, which, due to the local proposal updates of MALA, always stays within one mode. From this data, the mean and standard deviation of the observed mode were then estimated as $\overline{\phi}^{*}=0.79$ and $\sigma=0.06$. Consequently, we used as the proposal sampler a Gaussian mixture model with two modes centered at $\pm\overline{\phi}^{*}$ and of width $\sigma$.
Since we are again using a linear CV here, we use the same linear positions schedule $z(t_{k})$ and constant velocity schedule $v_{z}(t_{k})$ which we also used for the Gaussian tunnel. We compare the performance of the sampling algorithm over a grid of parameters $\alpha_{1}$, $\alpha_{2}$ and $\widetilde{v}$, using the mode jump cost as a criterion (the average number of calls to the force between transitions between $\overline{\phi}>0$ and $\overline{\phi}<0$). Optimal performance is again observed for the deterministic dynamics at $\alpha_{2}=0$ (see Fig.˜16 in Section˜H.3). Notably, as can be seen in Fig.˜4, there is a performance increase by around two orders of magnitude passing from the overdamped to the deterministic dynamics. This is consistent with the experiments for the Gaussian tunnel. Away from the optimum, we see a white diagonal where no transitions were observed, corresponding to a particularly ‘bad’ physical transition time $T/\sqrt{\beta M}=K_{T}\sqrt{\alpha_{2}}$. This might be a resonance effect due to the deterministic dynamics, but it is in any case far away from the parameter range of optimum performance.
For the optimal parameters, we also run a modified experiment where we deliberately bias the CV proposal sampler by putting a larger weight on the mode corresponding to negative magnetization. As we show in Fig.˜5, this bias is accurately corrected and the obtained samples align perfectly with the target density.
Driving the system from one mode to another corresponds to creating a domain of opposite orientation from either end of the lattice and gradually increasing its size. We indeed observe in Fig.˜6 that ‘good’ transitions with a high probability of acceptance follow one of these two transition paths. The existence of multiple transition paths does not pose a difficulty per se for the algorithm presented here: the acceptance of a proposal move is only determined by the work accumulated along the way and thus the intermediate bimodality is not a problem for accurate sampling of the two modes and their relative weights.
@@
Figure 5: Marginal distribution of the $\phi^{4}$ model obtained from running 8 chains for $10^{4}$ steps with $\alpha_{1}=0$, $\alpha_{2}=0.0014$ and $\widetilde{v}=1.2\times 10^{-3}$ (the optimal parameters obtained from optimization in Fig.˜4). In this case, a biased GMM proposal with weights $0.3$ and $0.7$ was used.
Figure 6: Two possible transition paths in $\phi^{4}$ model from positive to negative values of $\overline{\phi}$ with the same parameters as in Fig.˜5

### v3

@@ II.2.4 Constrained dynamics and work
where $({\mathcal{G}}^{k})$ and $({\mathcal{G}}^{k+1/2})$ are independent sequences of i.i.d. centered Gaussian random vectors with covariance matrix $\mathrm{Id}_{d}$. Note that the constraint in $(C_{q})$ is time-dependent, and thus the constraint in $(C_{p})$ is non zero, in contrast to the standard RATTLE scheme. The Lagrange multiplier $\lambda^{k+1/2}$ (resp. $\lambda^{k+3/4}$) is still chosen so that the constraint $(C_{q})$ (resp. $(C_{p})$) is satisfied. For $\lambda^{k+3/4}$, this leads to a linear equation with an analytical solution (as for the usual RATTLE scheme), whereas the solution for $\lambda^{k+1/2}$ will in general be obtained numerically using Newton’s methods. For further details on this point, in particular regarding uniqueness of the solution, see Remark˜B.3 in the Appendix. Note that there is an equivalent formulation of (II.17) and (II.19) that is more suitable for numerical implementation, which we provide in Appendix˜A for completeness. Besides, let us emphasize that the matrices $\gamma_{P}(q^{k})=P_{M}(q)\gamma P_{M}(q)^{\top}$ and $\sigma_{P}(q^{k})=P_{M}(q)\sigma$ satisfy the fluctuation-dissipation identity
@@
Assuming that the position and velocity schedules are compatible in the continuous-time limit $\Delta t\to 0$, i.e. $v_{z}(t)=\dot{z}(t)$, Eqs.˜II.17, II.18 and II.19 are discretizations of Langevin dynamics with constraints $(\xi(q(t)),v_{\xi}(q(t),p(t)))=(z(t),\dot{z}(t))$, see Ref. Lelièvre et al., 2012.
@@
for $k=0\ldots K_{T}-1$, where $\widetilde{H}$ denotes the Hamiltonian that is modified by the Fixman term (II.16), namely $\widetilde{H}(q,p)=\widetilde{V}(q)+\frac{1}{2}p^{\top}M^{-1}p$}.
For $\gamma=0$, the dynamics reduce to the deterministic part (II.18) (driven Hamiltonian dynamics) since (II.17) and (II.19) leave the system unchanged. The only remaining source of randomness is thus the initialization of the momentum $p^{0}$ in Step 2. Furthermore, since in this case $p^{k+1/4}=p^{k}$ and $p^{k+3/4}=p^{k+1}$, the work simplifies to a telescopic sum and just depends on the endpoints of the path:
@@
The idea of resampling the momentum at each iteration of the algorithm to build an MCMC sampler in position space is akin to Hamiltonian Monte-Carlo (HMC) Duane et al. (1987) and for $\gamma=0$, this analogy is particularly apparent. Unlike HMC, however, where the total energy is conserved up to a time discretization error, our algorithm does not conserve the energy due to the time-evolving constraints in CV space. Rather, the accumulated work is used in Step 3 of the algorithm to guarantee the reversibility of the Markov kernel.
Remark II.2.
When the Gram tensor $G_{M}(q)$ does not depend on $q$ (an example is given in Section˜III.3), the Fixman term just adds an overall shift to the energy $V$ and need not be considered. In this case, the requirement (II.6) on the initial and final velocity can also be relaxed to $v_{z}(t_{0})=v_{z}(t_{K_{T}})$ and can be non-zero, assuming that the initial momentum $p^{0}$ is still sampled from the appropriate subspace (see Remarks˜B.2 and B.10). This is true in particular for the case of a linear CV, which, after a change of basis, can always be seen as representing a subset of $\ell$ degrees of freedom. This is the standard setting treated in previous worksAthènes (2002); Nilmeier et al. (2011); Chen and Roux (2015a, b); Neal (2005), where some degrees of freedom are driven through a non-equilibrium protocol. We spell out the dynamics for the simplified setting of the linear CV in Appendix˜D, since it allows for an intuitive understanding as an alchemical transition with the CV as the alchemical parameter.
The problem of constructing a steered path can be understood as transporting samples from the constrained target measure at the initial CV value $Z_{n}$ to the final one $\widetilde{Z}_{n+1}$. Our method is based on a stochastic differential equation (with the auxiliary variable of the momentum), but other dynamics are possible, notably deterministic transformationsChen and Roux (2015a) or a combination of stochastic and deterministic steps, where the underlying SDE is modified by a deterministic ‘escorting’ termVaikuntanathan and Jarzynski (2011). The latter can also be understood as an instance of a stochastic normalizing flowCaselle et al. (2022). These approaches are not easy to adapt to our context of a nonlinear CV, however, since they require to transport samples supported on the submanifolds $\Sigma_{z}$.
@@ II.3 Reversibility of the algorithm
We prove this statement in Appendix˜B. The basic ingredient of the proof is to introduce a backward version of the dynamics (associated to the forward dynamics) and use a version of the Jarzynski–Crooks equalityCrooks (1999).
As a consequence of Theorem˜II.3, the MCMC algorithm admits $\nu$ as an invariant measure and, up to checking irreducibility which has to be done on a case-by-case basis, empirical averages over trajectories $(Q_{n})_{n\geq 0}$ converge to averages with respect to $\nu$.
Remark II.4.
@@ II.4 A normalized version of the algorithm
The overall algorithm depends on many parameters: $\Delta t$, $\gamma$, $\sigma$, $M$, the choice of the schedule $z(t)$ and $v_{z}(t)$, and the transition time $T$. In order to reduce the number of parameters, we will use the following normalized version in the numerical examples: Assuming that the mass matrix $M=M\,\mathrm{Id}$ is isotropic, we show in Appendix˜C that the overall dynamics only depends on the two parameters
@@
Regarding the physical interpretation of $\alpha_{1}$ and $\alpha_{2}$, note first that the damping term in (II.17) and (II.19) is associated to the decorrelation of momenta over a time scale of $M/\gamma$. The parameter $\alpha_{1}$ therefore measures the inverse of this friction decorrelation time, rescaled by the time discretization. Furthermore, the equilibrium distribution of the velocities is a centered normal distribution with covariance $(\beta M)^{-1}$, which means that $\alpha_{2}$ is the squared displacement generated by a typical thermal velocity within one integration step.
For the parameter choice $\alpha_{1}=1$, the steered dynamics (II.17)-(II.19) reduce to overdamped Langevin dynamics (see Appendix˜D). If $\xi$ is linear in addition, so that after a suitable change of basis the CV can be identified with the first $\ell$ degrees of freedom, the overall algorithm is equivalent to the ‘asymmetric’ algorithm presented in our earlier work Schönle et al. (2025) (up to a slight change of the definition of the steered path: here we first move the CV and then the remaining degrees of freedom). Therein, we also provided the continuous-time limit (which is the one of infinite damping or zero mass) for this version of the algorithm together with the corresponding stochastic differential equation (SDE). For the choice of $\alpha_{1}=0$, the steered dynamics correspond to the Verlet scheme (Leapfrog dynamics) in the orthogonal degrees of freedom and are therefore a discretization of Hamiltonian dynamics with a time-dependent Hamiltonian. Note that in this case, the continuous-time limit of the dynamics $\Delta t\to 0$ simply corresponds to taking $\alpha_{2}\to 0$ and the acceptance of the move in Step 3 is governed only by the (normalized) physical transition time $T/\sqrt{\beta M}=K_{T}\Delta t/\sqrt{\beta M}=\mathrm{ceil}\left(|\widetilde{Z}_{n+1}-Z_{n}|/\widetilde{v}\right)\sqrt{\alpha_{2}}$ (see Fig.˜15 in the appendix for a numerical illustration).
@@ III Numerical examples
In this section, we apply the proposed CV-guided sampler to examples of increasing complexity, demonstrating in particular the advantage of deterministic underdamped steering dynamics compared to the previously considered overdamped version. We consider four model systems of increasing complexity: a theoretical model-system dubbed ‘Gaussian tunnel’ (Section˜III.1), the one-dimensional $\phi^{4}$ model from statistical physics (Section˜III.2), a dimer in a solvent (Section˜III.3), and finally a 9-bead polymer in a solvent (Section˜III.4). For each example, we performed a grid search within our chosen parameterization to optimize the performance of the algorithm. For the more realistic polymer example, we consider a 27-dimensional CV for which a generative model is necessary to build a proposal sampler adapted to the high-dimensional CV space. In our experiments, we demonstrate the efficacy of our approach using a normalizing flow to propose moves in CV space and converging the Monte Carlo simulation on the fully solvated system.
@@ III.1 Gaussian tunnel
We first consider a simple model with a linear CV as already introduced in previous workSchönle et al. (2025). It is a toy model in $d=20$ dimensions with a designated one-dimensional variable $z$ serving as the CV and d1d-1 additional coordinates $(x^{\perp}_{i})_{1\leq i\leq d-1}$. The distribution of $z$ is given by a Gaussian mixture and the conditional distribution of the remaining coordinates is a Gaussian with a mean depending on $z$. More precisely,
@@
Running a rough optimization through a grid search reveals that the optimum performance is obtained for $\alpha_{1}=0$, which corresponds to a deterministic steered dynamics (see Fig.˜14 of Section˜H.1). We compare the deterministic ($\alpha_{1}=0$) and the overdamped ($\alpha_{1}=1$) steered dynamics in Fig.˜2 using the two performance measures, which show consistent results. The deterministic dynamics offer a large performance increase over the overdamped dynamics with a difference of around two orders of magnitude. In both cases, we observed that for too large values of $\alpha_{2}$ (a too coarse time discretization), the dynamics become unstable and the trajectories diverge, leading to zero acceptance of the overall move. The optimal value for $\alpha_{2}$ is then just below this threshold. Typical driven transition paths are shown for the optimum of the overdamped and the deterministic case in Fig.˜3. To verify the unbiasedness of the algorithm at the optimum, we show the marginals of $z$ and $x_{0}^{\perp}$ in Fig.˜4. They indeed agree with the target measure and show that the algorithm corrects for the biased proposal distribution in CV space.
@@ III.2 $\phi^{4}$ Model
As a second example we consider a field model from statistical physics, the one-dimensional $\phi^{4}$ model (also considered previouslySchönle et al. (2025)). It consists of $N$ degrees of freedom $\{\phi_{i}\}_{i=1}^{N}$ on a one-dimensional lattice and can be seen as a continuous version of the Ising model for ferromagnetism in solids. The target distribution is the Boltzmann–Gibbs probability distribution $\nu({\rm d}\phi)$ with the energy (using the simplified version of the model considered in Ref. Gabrié et al., 2022)
@@
We use the magnetization $\overline{\phi}$ as the collective variable, which is a natural order parameter of the system. We run the algorithm on the system in a wavelet basis (see Appendix˜E for details) since this order parameter then naturally emerges as the first component of the system. To fix the proposal sampler in CV space, we first ran a short simulation using the Metropolis Adjusted Langevin Algorithm (MALA) Rossky et al. (1978); Roberts and Tweedie (1996), which, due to the local proposal updates of MALA, always stays within one mode. From this data, the mean and standard deviation of the observed mode were then estimated as $\overline{\phi}^{*}=0.79$ and $\sigma=0.06$. Consequently, we used as the proposal sampler a Gaussian mixture model with two modes centered at $\pm\overline{\phi}^{*}$ and of width $\sigma$.
Since we are again using a linear CV here, we use the same linear positions schedule $z(t_{k})$ and constant velocity schedule $v_{z}(t_{k})$ which we also used for the Gaussian tunnel. We compare the performance of the sampling algorithm over a grid of parameters $\alpha_{1}$, $\alpha_{2}$ and $\widetilde{v}$, using the mode jump cost as a criterion (the average number of calls to the force between transitions between $\overline{\phi}>0$ and $\overline{\phi}<0$). Optimal performance is again observed for the deterministic dynamics at $\alpha_{2}=0$ (see Fig.˜17 in Section˜H.3). Notably, as can be seen in Fig.˜5, there is a performance increase by around two orders of magnitude passing from the overdamped to the deterministic dynamics. This is consistent with the experiments for the Gaussian tunnel. Away from the optimum, we see a white diagonal where no transitions were observed, corresponding to a particularly ‘bad’ physical transition time $T/\sqrt{\beta M}=K_{T}\sqrt{\alpha_{2}}$. This might be a resonance effect due to the deterministic dynamics, but it is in any case far away from the parameter range of optimum performance.
For the optimal parameters, we also run a modified experiment where we deliberately bias the CV proposal sampler by putting a larger weight on the mode corresponding to negative magnetization. As we show in Fig.˜6, this bias is accurately corrected and the obtained samples align perfectly with the target density.
Driving the system from one mode to another corresponds to creating a domain of opposite orientation from either end of the lattice and gradually increasing its size. We indeed observe in Fig.˜7 that ‘good’ transitions with a high probability of acceptance follow one of these two transition paths. The existence of multiple transition paths does not pose a difficulty per se for the algorithm presented here: the acceptance of a proposal move is only determined by the work accumulated along the way and thus the intermediate bimodality is not a problem for accurate sampling of the two modes and their relative weights.
@@
Figure 6: Marginal distribution of the $\phi^{4}$ model obtained from running 8 chains for $10^{4}$ steps with $\alpha_{1}=0$, $\alpha_{2}=0.0014$ and $\widetilde{v}=1.2\times 10^{-3}$ (the optimal parameters obtained from optimization in Fig.˜5). In this case, a biased GMM proposal with weights $0.3$ and $0.7$ was used.
Figure 7: Two possible transition paths in $\phi^{4}$ model from positive to negative values of $\overline{\phi}$ with the same parameters as in Fig.˜6